OpenStudy (anonymous):

y-2/y+9=1/6 not sure how to solve.

OpenStudy (anonymous):

Try multiplying both sides of the equation by y.

OpenStudy (anonymous):

by y-2 or by y+9?

OpenStudy (saifoo.khan):

\[\frac{y-2}{y+9}=\frac16\]does that look like that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Simple way, cross multiply, \[6(y-2)=1(y+9)\]

OpenStudy (anonymous):

thank you :}

OpenStudy (saifoo.khan):

No problem!